OpenStudy (anonymous):

In the binary number system which is used in computer operations, there are only two digits allowed: 0 and 1. A binary code is a system of binary numbers with a fixed number of digits that are used to represent letters, numbers, and symbols. To produce enough binary numbers to represent all of the letters of our alphabet (both upper and lower case), how many binary digits must be used?

OpenStudy (freckles):

There are 26 upper case letters and 26 lower case letters in the alphabet wouldn't the sum of those be enough binary numbers...?

OpenStudy (freckles):

like i could be wrong and not know enough about binary numbers

OpenStudy (freckles):

say we wanted to represent 1,2,3 are binary numbers they would all 3 be different binary numbers but there would only be 3 because we only need to represent 3 digits

OpenStudy (freckles):

for example 1=01 2=10 3=11

OpenStudy (freckles):

we have 3 different binary numbers for the three different digits we wanted to represent as binary numbers

OpenStudy (anonymous):

it says at the back of the book that the answer is "at least 6".. and I just wanted to know why 6?

OpenStudy (freckles):

oh! i understand what they are asking

OpenStudy (freckles):

like 1 through 3 can be represented by two digits like 1=01 2=10 3=11 we needed 2 binary digits

OpenStudy (freckles):

so we need 26+26=54 i wonder what we would get if we tried to represent 54 in binary

OpenStudy (anonymous):

ohkay? lol wait.. i'm really confused O.o

OpenStudy (freckles):

pretend we took the numbers 1 through 3 and we converted them to binary

OpenStudy (freckles):

we only need two base two place values 2^0 and 2^1

OpenStudy (anonymous):

okkkaayyy?

OpenStudy (freckles):

do you understand that part?

OpenStudy (anonymous):

why 2^0?

OpenStudy (freckles):

binary is based two

OpenStudy (freckles):

\[1_{10}=0 \cdot 2^1+1 \cdot 2^0=01_2 \\ 2_{10}=1 \cdot 2^1 +0\cdot 2^0=10_2 \\ 3_{10}=1 \cdot 2^1+1\cdot 2^0=11_2\]

OpenStudy (freckles):

so all you have to do is take the highest number in your set to figure out how many binary digits you will need for that set

OpenStudy (freckles):

take the highest number in your set and write it in binary

OpenStudy (freckles):

so 54 is the highest in our set

OpenStudy (freckles):

convert that number to binary

OpenStudy (freckles):

jee i don't know how to add 26+26=52

OpenStudy (anonymous):

26+26=52 not 54 though? O.o

OpenStudy (freckles):

so 52 is the highest convert that to binary

OpenStudy (freckles):

it was a type-o

OpenStudy (freckles):

you can add for me lol

OpenStudy (anonymous):

lol you type too fast.. can you do it slowly?

OpenStudy (anonymous):

this is what i understand so far .. you said let's pretend that we took numbers 1, 2, &3 so since only 1 & 0 can be used as binary numbers..... we have to...

OpenStudy (anonymous):

brb, dinner..

OpenStudy (freckles):

when we converted 3 to binary number 11 we seen that we only needed two digits for each number from 1 to 3 if we had {1,2,3,4,5,6,7,8,9,10} 10=1*2^3+0*2^2+1*2^1+0^2^0 so we only need 4 binary place values the one's digit, the two's digit, the four's digit, the eight's digit

OpenStudy (freckles):